OpenStudy (anonymous):

a swimmer of mass 55 kg jump from the edge of a cliff 30 m high. Touching the water 3 s after launch, what is the value of the frictional force exerted by the air on the swimmer?

OpenStudy (anonymous):

let first just get the acceleration which the body have to reach the pool. \[x =0.5 a t ^{2}\] \[a= 2x /t ^{2} = 2* 30 /9 = 6.6667\] Now the Total force is the result of subtraction between the resistance and gravity because  so the resultant force is the follwoing \[F total = Fg - Fresis.\] \[ma = mg - Fresis\] \[Fresis = m(g-a) = 55 (9.8 - 20/3) = 172.333\approx 172.3 N.\]